OpenStudy (anonymous):

A copy machine makes 24 copies per minute. How many copies does it make in 4 minutes and 30 seconds?

OpenStudy (anonymous):

108 copies. Rate = 24/1 Time= 4 + 30 sec = 4 + 1/2 (as 30 sec = 1/2 min) = 9/2 Now, Copies = R x T = 24 x 9/2 =108

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

you can either set up a proportion or figure it out just by reading the problem carefully. 24 copies per minute... 4 min 30 sec. 24 x 4 = 96 30 sec = half a minute, so take half of 24, which is 12, and add it to 96 and you get 108 copies in 4 hrs and 30 sec.

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

no problem :)